# Random Feature Representation Boosting

### v3

@@ E.1 SWIM Random Features
where $\epsilon>0$ and $f$ is the true target function, i.e., the mapping $x_{i}$ to the true label $y_{i}=f(x_{i})$. Specifically, to avoid quadratic time complexity in dataset size $n$, we first consider the $n$ points $x^{(1)}_{i}$, $i=1,...,n$, and then uniformly generate $n$ offset points $x^{(2)}_{i}=x^{(1)}_{i+j_{i}}$. We then sample $H$ points from these proportional to $q(\cdot,\cdot|\Phi)$. This ensures the sampling procedure remains linear with respect to dataset size $n$.
@@ E.2 Model Architectures and Hyperparameter Ranges for OpenML Tasks
Dataset ID Logistic Regression RFRBoost RFNN E2E MLP ResNet XGBoost
3 0.974 0.996 0.989 0.996 0.997
6 0.768 0.912 0.890 0.944 0.903
11 0.878 0.981 0.979 0.934 0.931
14 0.812 0.844 0.804 0.828 0.830
15 0.964 0.966 0.963 0.964 0.960
16 0.948 0.974 0.945 0.973 0.954
18 0.728 0.739 0.727 0.739 0.721
22 0.815 0.837 0.812 0.840 0.795
23 0.512 0.551 0.553 0.532 0.547
28 0.970 0.991 0.980 0.989 0.978
29 0.854 0.854 0.845 0.852 0.862
31 0.761 0.762 0.756 0.743 0.744
32 0.946 0.993 0.990 0.992 0.986
37 0.762 0.758 0.760 0.767 0.762
38 0.968 0.979 0.977 0.977 0.988
44 0.931 0.950 0.936 0.941 0.952
50 0.983 0.983 0.980 0.976 0.995
54 0.801 0.849 0.829 0.843 0.772
151 0.760 0.796 0.784 0.792 0.847
182 0.861 0.909 0.908 0.911 0.913
188 0.641 0.664 0.643 0.635 0.652
307 0.838 0.993 0.972 0.989 0.924
458 0.994 0.994 0.993 0.988 0.990
469 0.187 0.188 0.196 0.193 0.178
1049 0.908 0.910 0.904 0.907 0.919
1050 0.894 0.889 0.898 0.892 0.890
1053 0.808 0.807 0.806 0.810 0.809
1063 0.841 0.845 0.833 0.841 0.839
@@
Dataset ID Logistic Regression RFRBoost RFNN E2E MLP ResNet XGBoost
1067 0.853 0.855 0.858 0.855 0.853
1068 0.925 0.927 0.928 0.928 0.935
1461 0.902 0.902 0.904 0.900 0.903
1462 0.988 1.000 1.000 1.000 0.997
1464 0.766 0.791 0.794 0.769 0.783
1475 0.478 0.545 0.534 0.570 0.601
1480 0.715 0.703 0.703 0.671 0.684
1486 0.943 0.949 0.946 0.950 0.959
1487 0.935 0.945 0.945 0.944 0.940
1489 0.748 0.876 0.864 0.906 0.903
1494 0.873 0.882 0.868 0.878 0.873
1497 0.701 0.929 0.907 0.944 0.998
1510 0.977 0.974 0.974 0.961 0.961
1590 0.850 0.858 0.856 0.851 0.865
4534 0.938 0.952 0.949 0.960 0.956
4538 0.478 0.554 0.513 0.606 0.660
6332 0.735 0.791 0.763 0.791 0.796
23381 0.626 0.640 0.624 0.566 0.612
23517 0.505 0.516 0.502 0.489 0.500
40499 0.997 0.997 0.992 0.999 0.985
40668 0.752 0.775 0.759 0.777 0.800
40701 0.867 0.936 0.923 0.951 0.953
40966 0.994 0.999 0.989 0.997 0.984
40975 0.932 0.997 0.977 0.999 0.994
40982 0.718 0.752 0.758 0.755 0.800
40983 0.971 0.986 0.987 0.988 0.983
40984 0.901 0.934 0.930 0.927 0.929
40994 0.961 0.957 0.957 0.946 0.944
41027 0.685 0.812 0.798 0.844 0.838
@@
Dataset ID Ridge Regression Greedy RFRBoost $A_{scalar}$ Greedy RFRBoost $A_{diag}$ Greedy RFRBoost $A_{dense}$ Gradient RFRBoost RFNN E2E MLP ResNet XGBoost
41021 0.233 0.233 0.232 0.231 0.231 0.232 0.265 0.249
44956 0.647 0.631 0.621 0.615 0.616 0.622 0.614 0.633
44957 0.688 0.362 0.246 0.238 0.233 0.347 0.316 0.197
44958 0.595 0.239 0.205 0.193 0.191 0.246 0.246 0.039
44959 0.589 0.304 0.275 0.277 0.286 0.305 0.311 0.245
44960 0.290 0.069 0.050 0.049 0.049 0.088 0.200 0.037
44962 0.446 0.448 0.448 0.446 0.447 0.447 0.474 0.452
44963 0.841 0.784 0.773 0.749 0.752 0.783 0.703 0.694
44964 0.521 0.443 0.406 0.395 0.398 0.443 0.369 0.340
44965 0.883 0.867 0.873 0.860 0.858 0.871 0.864 0.839
44966 0.719 0.724 0.718 0.719 0.721 0.726 0.731 0.720
44967 0.786 0.788 0.790 0.788 0.784 0.795 0.814 0.789
44969 0.409 0.084 0.027 0.022 0.025 0.061 0.058 0.086
44970 0.645 0.609 0.610 0.618 0.615 0.603 0.616 0.622
44971 0.840 0.789 0.779 0.780 0.777 0.785 0.729 0.685
44972 0.796 0.776 0.775 0.768 0.767 0.774 0.777 0.722
44973 0.602 0.287 0.235 0.194 0.194 0.284 0.177 0.244
44974 0.457 0.214 0.184 0.153 0.152 0.211 0.158 0.126
44975 0.026 0.027 0.026 0.026 0.026 0.026 0.056 0.109
44976 0.270 0.192 0.177 0.170 0.170 0.196 0.161 0.192
44977 0.590 0.525 0.517 0.508 0.506 0.525 0.486 0.446
44978 0.242 0.149 0.145 0.146 0.148 0.149 0.142 0.134
44979 0.252 0.163 0.148 0.142 0.140 0.165 0.145 0.142
44980 0.766 0.430 0.324 0.318 0.317 0.438 0.275 0.457
44981 0.914 0.913 0.913 0.912 0.913 0.914 0.636 0.611
44983 0.431 0.263 0.250 0.247 0.244 0.257 0.235 0.229
44984 0.689 0.660 0.658 0.656 0.657 0.659 0.656 0.656
44987 0.334 0.233 0.194 0.183 0.192 0.246 0.229 0.211
44989 0.308 0.293 0.287 0.276 0.274 0.295 0.304 0.263
44990 0.404 0.395 0.396 0.395 0.393 0.399 0.405 0.399
44993 0.799 0.776 0.773 0.770 0.769 0.775 0.787 0.773
44994 0.265 0.215 0.216 0.215 0.216 0.216 0.231 0.225
45012 0.448 0.343 0.339 0.330 0.330 0.341 0.351 0.331
45402 0.621 0.526 0.507 0.499 0.497 0.532 0.474 0.502

### v4

@@ E.1 SWIM Random Features
where $\epsilon>0$ and $f$ is the true target function, i.e., the mapping $x_{i}$ to the true label $y_{i}=f(x_{i})$. Specifically, to avoid quadratic time complexity in dataset size $n$, we first consider the $n$ points $x^{(1)}_{i}$, i=1,,ni=1,...,n, and then uniformly generate $n$ offset points $x^{(2)}_{i}=x^{(1)}_{i+j_{i}}$. We then sample $H$ points from these proportional to $q(\cdot,\cdot|\Phi)$. This ensures the sampling procedure remains linear with respect to dataset size $n$.
@@ E.2 Model Architectures and Hyperparameter Ranges for OpenML Tasks
Dataset
ID
Logistic
Regression
RFRBoost
RFNN
E2E MLP
ResNet
XGBoost
3 0.974 (0.005) 0.996 (0.002) 0.989 (0.004) 0.996 (0.001) 0.997 (0.002)
6 0.768 (0.005) 0.912 (0.008) 0.890 (0.009) 0.944 (0.009) 0.903 (0.015)
11 0.878 (0.033) 0.981 (0.006) 0.979 (0.004) 0.934 (0.042) 0.931 (0.031)
14 0.812 (0.011) 0.844 (0.017) 0.804 (0.018) 0.828 (0.020) 0.830 (0.008)
15 0.964 (0.016) 0.966 (0.011) 0.963 (0.017) 0.964 (0.015) 0.960 (0.012)
16 0.948 (0.004) 0.974 (0.009) 0.945 (0.008) 0.973 (0.007) 0.954 (0.016)
18 0.728 (0.010) 0.739 (0.010) 0.727 (0.010) 0.739 (0.007) 0.721 (0.017)
22 0.815 (0.012) 0.837 (0.004) 0.812 (0.011) 0.840 (0.017) 0.795 (0.012)
23 0.512 (0.054) 0.551 (0.038) 0.553 (0.044) 0.532 (0.035) 0.547 (0.022)
28 0.970 (0.005) 0.991 (0.003) 0.980 (0.006) 0.989 (0.005) 0.978 (0.004)
29 0.854 (0.017) 0.854 (0.024) 0.845 (0.024) 0.852 (0.027) 0.862 (0.026)
31 0.761 (0.023) 0.762 (0.028) 0.756 (0.028) 0.743 (0.039) 0.744 (0.031)
32 0.946 (0.002) 0.993 (0.003) 0.990 (0.002) 0.992 (0.001) 0.986 (0.003)
37 0.762 (0.020) 0.758 (0.027) 0.760 (0.032) 0.767 (0.025) 0.762 (0.034)
38 0.968 (0.005) 0.979 (0.005) 0.977 (0.004) 0.977 (0.004) 0.988 (0.004)
44 0.931 (0.007) 0.950 (0.004) 0.936 (0.009) 0.941 (0.009) 0.952 (0.002)
50 0.983 (0.006) 0.983 (0.010) 0.980 (0.010) 0.976 (0.016) 0.995 (0.007)
54 0.801 (0.026) 0.849 (0.018) 0.829 (0.021) 0.843 (0.015) 0.772 (0.019)
151 0.760 (0.015) 0.796 (0.014) 0.784 (0.015) 0.792 (0.010) 0.847 (0.013)
182 0.861 (0.009) 0.909 (0.007) 0.908 (0.007) 0.911 (0.009) 0.913 (0.006)
188 0.641 (0.057) 0.664 (0.049) 0.643 (0.031) 0.635 (0.038) 0.652 (0.029)
307 0.838 (0.012) 0.993 (0.002) 0.972 (0.016) 0.989 (0.010) 0.924 (0.031)
458 0.994 (0.000) 0.994 (0.007) 0.993 (0.006) 0.988 (0.013) 0.990 (0.006)
469 0.187 (0.023) 0.188 (0.028) 0.196 (0.021) 0.193 (0.028) 0.178 (0.031)
1049 0.908 (0.012) 0.910 (0.013) 0.904 (0.008) 0.907 (0.014) 0.919 (0.009)
1050 0.894 (0.013) 0.889 (0.010) 0.898 (0.010) 0.892 (0.006) 0.890 (0.007)
1053 0.808 (0.009) 0.807 (0.011) 0.806 (0.012) 0.810 (0.011) 0.809 (0.014)
1063 0.841 (0.023) 0.845 (0.039) 0.833 (0.024) 0.841 (0.020) 0.839 (0.021)
@@
Dataset
ID
Logistic
Regression
RFRBoost
RFNN
E2E MLP
ResNet
XGBoost
1067 0.853 (0.019) 0.855 (0.018) 0.858 (0.015) 0.855 (0.016) 0.853 (0.012)
1068 0.925 (0.008) 0.927 (0.016) 0.928 (0.009) 0.928 (0.010) 0.935 (0.012)
1461 0.902 (0.008) 0.902 (0.005) 0.904 (0.009) 0.900 (0.006) 0.903 (0.005)
1462 0.988 (0.005) 1.000 (0.000) 1.000 (0.000) 1.000 (0.000) 0.997 (0.004)
1464 0.766 (0.039) 0.791 (0.037) 0.794 (0.046) 0.769 (0.048) 0.783 (0.043)
1475 0.478 (0.005) 0.545 (0.010) 0.534 (0.012) 0.570 (0.005) 0.601 (0.008)
1480 0.715 (0.037) 0.703 (0.017) 0.703 (0.011) 0.671 (0.033) 0.684 (0.020)
1486 0.943 (0.006) 0.949 (0.003) 0.946 (0.004) 0.950 (0.006) 0.959 (0.003)
1487 0.935 (0.012) 0.945 (0.012) 0.945 (0.012) 0.944 (0.009) 0.940 (0.011)
1489 0.748 (0.010) 0.876 (0.015) 0.864 (0.005) 0.906 (0.005) 0.903 (0.006)
1494 0.873 (0.016) 0.882 (0.016) 0.868 (0.018) 0.878 (0.015) 0.873 (0.010)
1497 0.701 (0.009) 0.929 (0.005) 0.907 (0.008) 0.944 (0.007) 0.998 (0.001)
1510 0.977 (0.004) 0.974 (0.008) 0.974 (0.014) 0.961 (0.023) 0.961 (0.013)
1590 0.850 (0.005) 0.858 (0.010) 0.856 (0.010) 0.851 (0.009) 0.865 (0.013)
4534 0.938 (0.007) 0.952 (0.010) 0.949 (0.003) 0.960 (0.009) 0.956 (0.008)
4538 0.478 (0.014) 0.554 (0.017) 0.513 (0.014) 0.606 (0.014) 0.660 (0.017)
6332 0.735 (0.031) 0.791 (0.024) 0.763 (0.009) 0.791 (0.023) 0.796 (0.021)
23381 0.626 (0.040) 0.640 (0.030) 0.624 (0.034) 0.566 (0.048) 0.612 (0.041)
23517 0.505 (0.010) 0.516 (0.009) 0.502 (0.017) 0.489 (0.013) 0.500 (0.012)
40499 0.997 (0.001) 0.997 (0.001) 0.992 (0.000) 0.999 (0.001) 0.985 (0.004)
40668 0.752 (0.017) 0.775 (0.014) 0.759 (0.019) 0.777 (0.010) 0.800 (0.014)
40701 0.867 (0.003) 0.936 (0.002) 0.923 (0.004) 0.951 (0.009) 0.953 (0.003)
40966 0.994 (0.005) 0.999 (0.002) 0.989 (0.007) 0.997 (0.002) 0.984 (0.016)
40975 0.932 (0.007) 0.997 (0.004) 0.977 (0.006) 0.999 (0.001) 0.994 (0.004)
40982 0.718 (0.028) 0.752 (0.016) 0.758 (0.024) 0.755 (0.028) 0.800 (0.018)
40983 0.971 (0.003) 0.986 (0.002) 0.987 (0.003) 0.988 (0.004) 0.983 (0.003)
40984 0.901 (0.009) 0.934 (0.012) 0.930 (0.007) 0.927 (0.011) 0.929 (0.011)
40994 0.961 (0.018) 0.957 (0.019) 0.957 (0.019) 0.946 (0.015) 0.944 (0.021)
41027 0.685 (0.004) 0.812 (0.007) 0.798 (0.012) 0.844 (0.007) 0.838 (0.006)
@@
Dataset
ID
Ridge
Regression
Greedy
RFRBoost
$A_{scalar}$
Greedy
RFRBoost
$A_{diag}$
Greedy
RFRBoost
$A_{dense}$
Gradient
RFRBoost
RFNN
E2E
MLP
ResNet
XGBoost
41021 0.233 (0.012) 0.233 (0.013) 0.232 (0.013) 0.231 (0.012) 0.231 (0.011) 0.232 (0.012) 0.265 (0.011) 0.249 (0.011)
44956 0.647 (0.020) 0.631 (0.018) 0.621 (0.019) 0.615 (0.019) 0.616 (0.020) 0.622 (0.022) 0.614 (0.023) 0.633 (0.020)
44957 0.688 (0.033) 0.362 (0.026) 0.246 (0.017) 0.238 (0.011) 0.233 (0.013) 0.347 (0.021) 0.316 (0.026) 0.197 (0.015)
44958 0.595 (0.010) 0.239 (0.019) 0.205 (0.010) 0.193 (0.016) 0.191 (0.027) 0.246 (0.022) 0.246 (0.025) 0.039 (0.011)
44959 0.589 (0.026) 0.304 (0.007) 0.275 (0.013) 0.277 (0.012) 0.286 (0.018) 0.305 (0.015) 0.311 (0.011) 0.245 (0.026)
44960 0.290 (0.028) 0.069 (0.013) 0.050 (0.005) 0.049 (0.005) 0.049 (0.006) 0.088 (0.024) 0.200 (0.027) 0.037 (0.007)
44962 0.446 (0.063) 0.448 (0.062) 0.448 (0.062) 0.446 (0.063) 0.447 (0.062) 0.447 (0.063) 0.474 (0.069) 0.452 (0.067)
44963 0.841 (0.011) 0.784 (0.005) 0.773 (0.020) 0.749 (0.008) 0.752 (0.005) 0.783 (0.013) 0.703 (0.017) 0.694 (0.010)
44964 0.521 (0.017) 0.443 (0.018) 0.406 (0.013) 0.395 (0.012) 0.398 (0.009) 0.443 (0.014) 0.369 (0.005) 0.340 (0.010)
44965 0.883 (0.066) 0.867 (0.071) 0.873 (0.077) 0.860 (0.075) 0.858 (0.070) 0.871 (0.077) 0.864 (0.075) 0.839 (0.081)
44966 0.719 (0.082) 0.724 (0.085) 0.718 (0.084) 0.719 (0.082) 0.721 (0.080) 0.726 (0.075) 0.731 (0.079) 0.720 (0.080)
44967 0.786 (0.065) 0.788 (0.071) 0.790 (0.068) 0.788 (0.065) 0.784 (0.067) 0.795 (0.069) 0.814 (0.045) 0.789 (0.056)
44969 0.409 (0.011) 0.084 (0.027) 0.027 (0.007) 0.022 (0.002) 0.025 (0.002) 0.061 (0.022) 0.058 (0.015) 0.086 (0.003)
44970 0.645 (0.042) 0.609 (0.044) 0.610 (0.044) 0.618 (0.035) 0.615 (0.040) 0.603 (0.045) 0.616 (0.029) 0.622 (0.043)
44971 0.840 (0.028) 0.789 (0.018) 0.779 (0.022) 0.780 (0.017) 0.777 (0.021) 0.785 (0.024) 0.729 (0.009) 0.685 (0.015)
44972 0.796 (0.036) 0.776 (0.028) 0.775 (0.030) 0.768 (0.026) 0.767 (0.030) 0.774 (0.026) 0.777 (0.029) 0.722 (0.012)
44973 0.602 (0.007) 0.287 (0.013) 0.235 (0.008) 0.194 (0.007) 0.194 (0.004) 0.284 (0.016) 0.177 (0.007) 0.244 (0.008)
44974 0.457 (0.014) 0.214 (0.013) 0.184 (0.015) 0.153 (0.008) 0.152 (0.007) 0.211 (0.012) 0.158 (0.012) 0.126 (0.012)
44975 0.026 (0.008) 0.027 (0.008) 0.026 (0.008) 0.026 (0.008) 0.026 (0.008) 0.026 (0.008) 0.056 (0.012) 0.109 (0.003)
44976 0.270 (0.004) 0.192 (0.007) 0.177 (0.009) 0.170 (0.005) 0.170 (0.006) 0.196 (0.009) 0.161 (0.006) 0.192 (0.007)
44977 0.590 (0.021) 0.525 (0.019) 0.517 (0.016) 0.508 (0.019) 0.506 (0.021) 0.525 (0.018) 0.486 (0.020) 0.446 (0.022)
44978 0.242 (0.016) 0.149 (0.005) 0.145 (0.007) 0.146 (0.006) 0.148 (0.007) 0.149 (0.006) 0.142 (0.004) 0.134 (0.011)
44979 0.252 (0.015) 0.163 (0.015) 0.148 (0.012) 0.142 (0.010) 0.140 (0.010) 0.165 (0.013) 0.145 (0.011) 0.142 (0.011)
44980 0.766 (0.022) 0.430 (0.008) 0.324 (0.007) 0.318 (0.008) 0.317 (0.007) 0.438 (0.022) 0.275 (0.010) 0.457 (0.031)
44981 0.914 (0.017) 0.913 (0.016) 0.913 (0.017) 0.912 (0.017) 0.913 (0.017) 0.914 (0.017) 0.636 (0.015) 0.611 (0.006)
44983 0.431 (0.004) 0.263 (0.012) 0.250 (0.013) 0.247 (0.016) 0.244 (0.014) 0.257 (0.010) 0.235 (0.019) 0.229 (0.014)
44984 0.689 (0.023) 0.660 (0.023) 0.658 (0.024) 0.656 (0.022) 0.657 (0.023) 0.659 (0.023) 0.656 (0.023) 0.656 (0.023)
44987 0.334 (0.035) 0.233 (0.020) 0.194 (0.024) 0.183 (0.026) 0.192 (0.036) 0.246 (0.019) 0.229 (0.057) 0.211 (0.020)
44989 0.308 (0.010) 0.293 (0.009) 0.287 (0.006) 0.276 (0.012) 0.274 (0.013) 0.295 (0.009) 0.304 (0.016) 0.263 (0.005)
44990 0.404 (0.009) 0.395 (0.009) 0.396 (0.008) 0.395 (0.009) 0.393 (0.009) 0.399 (0.010) 0.405 (0.008) 0.399 (0.008)
44993 0.799 (0.017) 0.776 (0.020) 0.773 (0.017) 0.770 (0.018) 0.769 (0.019) 0.775 (0.017) 0.787 (0.018) 0.773 (0.019)
44994 0.265 (0.016) 0.215 (0.011) 0.216 (0.009) 0.215 (0.007) 0.216 (0.008) 0.216 (0.009) 0.231 (0.013) 0.225 (0.010)
45012 0.448 (0.027) 0.343 (0.020) 0.339 (0.019) 0.330 (0.018) 0.330 (0.017) 0.341 (0.016) 0.351 (0.022) 0.331 (0.021)
45402 0.621 (0.023) 0.526 (0.024) 0.507 (0.017) 0.499 (0.017) 0.497 (0.019) 0.532 (0.019) 0.474 (0.030) 0.502 (0.029)